# A Survey on Monocular Re-Localization: From the Perspective of Scene Map Representation

### v1

@@ 8.3.2 MRL  for  FPE
VLAD tokyo247 0.21,12.5° 0.33,13.8° 0.15,14.9° 0.28,11.2° 0.31,11.2° 0.30,11.3° 0.25,12.3° 0.26,12.46°
VLAD+InterSattler2019UnderstandingTL 0.18,10.0° 0.33,12.4° 0.14,14.3° 0.25,10.1° 0.26,9.4° 0.27,11.1° 0.24,14.7° 0.24,11.71°
@@
NN-Net Zakaria2017ICCVW 0.13,6.5° 0.26,12.7° 0.14,12.3° 0.21,7.4° 0.24,6.4° 0.24,8.0° 0.27,11.8° 0.21,9.30°
RelocNetRelocNet 0.12,4.1° 0.26,10.4° 0.14,10.5° 0.18,5.3° 0.26,4.2° 0.23,5.1° 0.28,7.5° 0.21,6.73°
EssNet EssNet 0.13,5.1° 0.27,10.1° 0.15,9.9° 0.21,6.9° 0.22,6.1° 0.23,6.9° 0.32,11.2° 0.22,8.03°
NC-EssNet EssNet 0.12,5.6° 0.26,9.6° 0.14,10.7° 0.20,6.7° 0.22,5.7° 0.22,6.3° 0.31,7.9° 0.21,7.50°
RelocGNNDSAC2 0.08,2.7° 0.21,7.5° 0.13,8.70° 0.15,4.1° 0.15,3.5° 0.19,3.7° 0.22,6.5° 0.16,5.24°
CamNet CamNet 0.04,1.73° 0.03, 1.74° 0.05,1.98° 0.04,1.62° 0.04,1.64° 0.04,1.63° 0.04,1.51° 0.04,1.69°
@@
Active Search VPS 0.04,2.0° 0.03,1.5° 0.02,1.5° 0.09,3.6° 0.08,3.1° 0.07,3.4 ° 0.03,2.2° 0.05,2.47°
InLocInLoc 0.03,1.05° 0.03,1.07° 0.02,1.16° 0.03,1.05° 0.05,1.55° 0.04,1.31° 0.09,2.47° 0.04,1.44°
@@
PoseNet posenet 0.32,8.12° 0.47,14.4° 0.29,12.0° 0.48,7.68° 0.47,8.42° 0.59,8.64° 0.47,13.8° 0.44,10.44°
BayesianPN BayesianPN 0.37,7.24° 0.43,13.7° 0.31,12.0° 0.48,8.04° 0.61,7.08° 0.58,7.54° 0.48,13.1° 0.47,9.81°
LSTM-PN LSTMPN 0.24,5.77° 0.34,11.9° 0.21,13.7° 0.30,8.08° 0.33,7.00° 0.37,8.83° 0.40,13.7° 0.31,9.85°
GPoseNet Cai2018AHP 0.20,7.11° 0.38,12.3° 0.21,13.8° 0.28,8.83° 0.37,6.94° 0.35,8.15° 0.37,12.5° 0.31,9.95°
PoseNetLearnPoseNet2 0.14,4.50° 0.27,11.8° 0.18,12.1° 0.20,5.77° 0.25,4.82° 0.24,5.52° 0.37,10.6° 0.24,7.87°
GeoPoseNetPoseNet2 0.13,4.48° 0.27,11.3° 0.17,13.0° 0.19,5.55° 0.26,4.75° 0.23,5.35 ° 0.35,12.4° 0.23,8.12°
IRPNetShavit2020DoWR 0.13,5.64° 0.25,9.67° 0.15,13.1° 0.24,6.33° 0.22,5.78° 0.30,7.29° 0.34,11.6° 0.23,8.49°
AtLocWang2019AtLocAG 0.10,4.07° 0.25,11.4° 0.16,11.8° 0.17,5.34° 0.21,4.37 ° 0.23,5.42° 0.26,10.5° 0.20,7.56°
@@
MSPNMSPN 0.09/4.76° 0.29/10.5° 0.16/13.1° 0.16/6.80° 0.19/5.50° 0.21/6.61° 0.31/11.6° 0.20/7.56°
MS-TransMS-Trans1 0.11,4.66° 0.24,9.60 ° 0.14,12.2° 0.17,5.66° 0.18,4.44° 0.17,5.94 ° 0.26,8.45° 0.18,7.28°
c2f-MS-Trans MS-Transformer2 0.10,4.60° 0.24,9.88° 0.12,12.3° 0.16,5.64° 0.16,4.42° 0.16,6.39° 0.25,7.76° 0.17/7.28°
@@
DSAC DSAC 0.02,0.7° 0.03,1.0° 0.02,1.30° 0.03,1.0° 0.05,1.30° 0.05,1.5° 1.90,49.4° 0.30,8.03°
DSAC++Less_is_More 0.02,0.5° 0.02,0.9° 0.01,0.8 0.03,0.7° 0.04,1.1° 0.04,1.1° 0.09,2.6° 0.04,1.10°
$*{}^{*}$ DSAC2 0.02,1.1° 0.02,1.0° 0.01,1.8° 0.03,1.2° 0.04,1.4° 0.03,1.7° 0.04,1.4° 0.03,1.37°
SANet SANet 0.03/0.9° 0.03/1.1° 0.02/1.5° 0.03/1.0° 0.05/1.3° 0.04/1.4° 0.16/4.6° n.a.
@@
DFNet DFNet 0.04,1.48° 0.04,2.16° 0.03,1.82° 0.07,2.01° 0.09,2.26° 0.09,2.42° 0.14,3.31° 0.07,2.21°
Direct-PN+U DPN 0.09,2.77° 0.16,4.87° 0.10,6.64° 0.17,5.04° 0.19,3.59° 0.19,4.79° 0.24,8.52° 0.16,5.17°
LENS Moreau2021LENSLE 0.03,1.3° 0.10,3.7° 0.07,5.8° 0.07,1.9° 0.08,2.2° 0.09,2.2° 0.14,3.6° 0.08,3.0°
@@
NeRF-Loc liu2023nerfloc 0.02/1.1° 0.02/1.1° 0.01/1.9° 0.02/1.1° 0.03/1.3° 0.03/1.5° 0.03/1.3° n.a.
NN-MRL: Then, we discuss about MRL with learnt implicit map in Sec. 7. According to the localization results on medium-scale Cambridge Landmarks dataset posenet in Tab. 8 and small-scale 7Scenes dataset posenet in Tab. 9, we find that the localization accuracy of APR methods is not pleased, and sometimes they even perform worse than VPR methods that only give an approximated pose. APR performs much worse than VL-MRL methods. When enhanced by NeRF, the performances of APR methods are significantly improved, and achieve comparable performances with VL-MRL methods. As a surprise, we find that SCR methods work well and achieve impressive localization accuracy in both medium-scale outdoor and small-scale indoor scenes. DSAC DSAC and its extended works Less_is_More ; DSAC2 even work better than VL-MRL methods in some cases. SCR may be a promising trend in MRL research in the future. But its heavy cost of training will block its wide usage. As for NeRF, it can significantly boost APR methods DFNet ; DPN ; Moreau2021LENSLE and it also can work well as a pose estimator liu2023nerfloc . We are still waiting for the usage of NeRF in large-scale and long-term localization tasks since building NeRF for unbounded scene is a much more challenging problem compared to indoor scene.
Summary: Finally, we compare many MRL methods with various kind of scene map on outdoor Cambridge Landmarks posenet and indoor 7Scenes posenet , as listed in Tab. 8 and Tab. 9. In outdoor scene, the rankings of MRL methods with regard to localization accuracy should be: NeRF pose estimator liu2023nerfloc $>$ I2P-MRL pixloc $>$ SCR $>$ APR+NeRF $\approx$ VL-MRL $>$ RPR $>$ APR $>$ VPR. And in indoor scene, the rankings of MRL methods are slightly changed as: NeRF pose estimator liu2023nerfloc $\approx$ I2P-MRL pixloc $\approx$ SCR $\approx$ APR+NeRF $\approx$ VL-MRL $>$ RPR $>$ APR $>$ VPR. It can be seen that in small-scale indoor scene, although APR and RPR work worse than other MRL methods, they can also achieve pleased localization performance. Given the low requirement of built map in APR and RPR, they can be used in some applications that do not need high-precision localization, such as VR. VPR only retrieve visually similar reference image and regard the reference pose as an approximated query pose, so its localization accuracy is very limited when reference images are few in outdoor scene. In indoor scene, most MRL methods with various kinds of scene map achieve comparable localization performance. But in outdoor scene, the gaps among different methods are more significant. Using NeRF as pose estimator like NeRF-Loc liu2023nerfloc , I2P-MRL pixloc , SCR, and APR+NeRF seem to be promising ways in MRL researches. Adapting NeRF-based localization to large-scale unbounded scenes, reducing the requirement of high-precision point cloud map for I2P-MRL, reducing the training cost of SCR should be explored to make these MRL methods more practical.
9 Opening Discussion
9.1 Are Explicit 3D Models Necessary in Visual Localization?
Traditional MRL methods require an explicit 3D scene map, e.g., visual landmarks map in VL-MRL, point cloud map in I2P-MRL, and vectorized map in HD-MRL. An explicit 3D model of the scene appears to be necessary in the early stage of MRL research. However, with the recent advancements in MRL, some techniques no longer require an explicit 3D model, such as VPR, RPR, and NN-MRL. Therefore, a natural question arises: are explicit 3D models necessary for accurate visual localization? This is an extension of a widely known question: are large-scale 3D models really necessary for accurate visual localization?3dmodel .
As shown in Tab. 8 and 9, it can be seen that some 2D image-based methods without explicit map, namely, VPR and RPR, perform worse than 3D model-based method with explicit map such as VL-MRL and I2P-MRL. However, 2D image-based methods with a learnt implicit map, i.e., NN-MRL, can achieve comparable or even better performances. Specially, NeRF-Loc liu2023nerfloc has been found to localize with centimeter-level accuracy, surpassing VL-MRL and I2P-MRL by a considerable margin. Thus, it is believed that the learnable implicit map has the potential to be a superior alternative to the explicit map for accurate localisation. The explicit 3D map is not strictly necessary in current deep learning era.
9.2 How to Select Specific Localization Method?
@@
As shown in the experimental results in Sec. 8.3, the utilized scene map of MRL usually determines the localization accuracy that a particular kind of MRL can achieve. 1) VPR methods estimate current pose by retrieving historical posed images. Their localization accuracy is limited, but they solely require a 2D image map and can obtain reference images with overlapped view-of-field. As a result, they are often utilized as a coarse step in hierarchical localization framework hloc1 ; hloc2 to acquire co-visible query-reference image pairs. 2) With a slightly better localization accuracy, RPR and APR methods can be employed in some applications that do not necessitate high-precision poses but call for a low-cost and light scene map. 3) VL-MRL and I2P-MRL methods require an offline mapping stage to obtain 3D scene map, i.e., visual landmarks map or point cloud map, so their cost of mapping is significantly higher than previously mentioned methods. With these precise and explicit 3D scene map, these two kind of methods achieve outstanding localization performance under most circumstances. However, when aiming for autonomy in large-scale scenarios, these two kinds of scene map require significant storage space, making them difficult to deploy on resource-limited platform, like intelligent vehicles. 4) HD map is the most light and compact form of scene map, so HD-MRL methods are widely applied in autonomous driving task. But HD Map is limited to human-made scenes with semantic map elements, and it is not suitable for natural scenes. 5) Recently developed NN-MRL methods utilize DNN model to represent scenes, which have demonstrated impressive success in small-scale indoor scene and medium-scale urban scenes. However, the efficacy of NN-MRL has yet to be evaluated in unbounded scenes. For instance, SCR methods require dense ground-truth data of scene coordinates, which is hard to be obtained in large-scale scenes. Training NeRF for unbounded dynamic scenes remains a formidable task, rendering MRL based on NeRF currently unsolvable.
@@
9.3 What’s Next for Visual Localization?
@@
Developed over serveral decades, MRL continues to attract growing attention from both academia and industry. We would like to discuss about an interesting and instructive problem: ” What are the future trends of visual localization research?”
@@
End-to-end pipeline: The robustness against environmental changes should be well considered by MRL methods for long-term autonomy. The current query image often has different visual conditions with reference data stored in the scene map. Most of existing solutions rely on image feature algorithms sift ; superpoint to attain condition-invariant image representation. However, these features often fail to achieve comparable performance to that on feature matching tasks. For instance, the matching accuracy of D2-Net d2net is inferior to SuperPoint superpoint on HPatches benchmark hpatches1 ; hpatches2 , but whtn it comes to localization benchmark benchmark6dof , D2-Net d2net yields better results than SuperPoint superpoint . This discrepancy between feature matching and localization might be attributed to the fact that human-defined supervision of features. Some recent approaches suggest building the whole localization pipeline into an E2E manner, incorporating feature extraction, matching, and pose estimation, and training such localization models by supervising them on estimated poses I2PNet ; I2DLoc ; Stumberg2019GNNetTG ; pixloc . Such pose supervision can facilitate the learning of involved geometry priors in MRL and enhance the global consistency of feature matching. This kind of image feature are specifically learned for MRL, enabling the closure of the gap between feature matching and localization, thus becoming a more reasonable solution for representation in MRL research. It is pertinent to note that HD-MRL is encountered with similar challenges alongside feature-based localization. Since semantic instance-level matching is sparse and difficult, some E2E solutions prefer to make models extract and match high-dimensional features for both images and vectorized map elements, and then estimate poses, leading to improved localization performances bevlocator ; he2023egovm . We believe E2E localization system will be a promising trend.
@@
Resource-friendly MRL: To achieve real-time ego-pose estimation in large-scale scenarios during long-term exploration, we must take into account the storage demands of scene map and the running efficiency of localization algorithms. For example, high-level autonomous vehicles in unbounded driving areas generally depend on vectorized HD map, rather than visual landmarks map and point cloud map that are commonly used in robotic applications within confined scenes roadsidemapping . In the field of robotic localization, several recommendations are available to lighten scene map, e.g., SceneSqueezer SceneSqueezer , NeuMap neumap , and HyperMap hypermap , which significantly decrease the map size while ensuring localization accuracy. There is no doubt that MRL will prefer a lightweight and compact scene map in the future. Aiming at high efficiency, MRL methods are commonly built in coarse-to-fine framework where VPR is applied to coarsely estimate poses hloc1 ; hloc2 . Recent advances in MRL methods enable us to employ NN-MRL solutions as a coarse step, such as APR in LATITUDE LATITUDE . Additionally, improving the efficiency of key components in practical MRL pipelines is also essential. ALIKE alike efficiently extracts local features while LightGLue lindenberger2023lightglue achieves high-speed matching. Efforts to reduce computational cost of MRL methods should be encouraged, as this will stimulate the proposal of a growing number of practical MRL solutions for resource-limited platforms.
@@
MRL with new kind of map: In recent years, numerous MRL methods that use new types of scene map have been created. For example, MeshLoc meshloc employs a dense mesh-based map rather than visual landmarks map and achieves superior performance. Similarly, NN-MRL methods utilize implicit neural network-based scene map as a substitute for traditional explicit 3D structure-based scene map, e.g., visual landmarks map and point cloud map. Proposing new representation formats of scene map is still an area of research that requires further exploration in MRL. The ideal scene map should be lightweight, compact, and easy to build and deploy. Moreover, it ought to provide comprehensive information for autonomy. The reference data for localization in the scene map should be sufficiently robust against potential changes in actual scene. Current scene map is challenging to accomplish all of these criteria, so we believe it is necessary to develop new kind of scene map for MRL, which will bring about impressive improvement to MRL researches.
@@
10 Conclusions
In this survey, we formulate MRL method as a interaction procedure between query image and scene map where poses are estimated, and then we systemically review MRL methods based on the representation format of utilized scene map, that is MRL using geo-tagged frame map, visual landmark map, point cloud map, HD map, and more recently proposed learnt implicit map. Each kind of MRL methods and the related component of them are fully reviewed. Also, we provide a review about evaluation of MRL methods, and draw some conclusions based on the evaluation results of typical algorithms. We provide some opening problem in this field and give our personal opinions. Finally, as a continuous contribution to the community, we list the reviewed papers and datasets on the website so that researches can readily find the best-matched MRL method based on their interests.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants U22A20104 and 52372414, and Beijing Municipal Science and Technology Commission (Grant No.Z221100008122011).

### v3

@@ IX-A Are Explicit 3D Models Necessary in Visual Localization?
VLAD [29] 0.21,12.5° 0.33,13.8° 0.15,14.9° 0.28,11.2° 0.31,11.2° 0.30,11.3° 0.25,12.3° 0.26,12.46°
VLAD+Inter[306] 0.18,10.0° 0.33,12.4° 0.14,14.3° 0.25,10.1° 0.26,9.4° 0.27,11.1° 0.24,14.7° 0.24,11.71°
@@
NN-Net [114] 0.13,6.5° 0.26,12.7° 0.14,12.3° 0.21,7.4° 0.24,6.4° 0.24,8.0° 0.27,11.8° 0.21,9.30°
RelocNet[22] 0.12,4.1° 0.26,10.4° 0.14,10.5° 0.18,5.3° 0.26,4.2° 0.23,5.1° 0.28,7.5° 0.21,6.73°
EssNet [107] 0.13,5.1° 0.27,10.1° 0.15,9.9° 0.21,6.9° 0.22,6.1° 0.23,6.9° 0.32,11.2° 0.22,8.03°
NC-EssNet [107] 0.12,5.6° 0.26,9.6° 0.14,10.7° 0.20,6.7° 0.22,5.7° 0.22,6.3° 0.31,7.9° 0.21,7.50°
@@
Active Search [151] 0.04,2.0° 0.03,1.5° 0.02,1.5° 0.09,3.6° 0.08,3.1° 0.07,3.4 ° 0.03,2.2° 0.05,2.47°
InLoc[35] 0.03,1.05° 0.03,1.07° 0.02,1.16° 0.03,1.05° 0.05,1.55° 0.04,1.31° 0.09,2.47° 0.04,1.44°
@@
PoseNet [27] 0.32,8.12° 0.47,14.4° 0.29,12.0° 0.48,7.68° 0.47,8.42° 0.59,8.64° 0.47,13.8° 0.44,10.44°
BayesianPN [241] 0.37,7.24° 0.43,13.7° 0.31,12.0° 0.48,8.04° 0.61,7.08° 0.58,7.54° 0.48,13.1° 0.47,9.81°
LSTM-PN [242] 0.24,5.77° 0.34,11.9° 0.21,13.7° 0.30,8.08° 0.33,7.00° 0.37,8.83° 0.40,13.7° 0.31,9.85°
AtLoc[246] 0.10,4.07° 0.25,11.4° 0.16,11.8° 0.17,5.34° 0.21,4.37 ° 0.23,5.42° 0.26,10.5° 0.20,7.56°
@@
MSPN[249] 0.09/4.76° 0.29/10.5° 0.16/13.1° 0.16/6.80° 0.19/5.50° 0.21/6.61° 0.31/11.6° 0.20/7.56°
MS-Trans[307] 0.11,4.66° 0.24,9.60 ° 0.14,12.2° 0.17,5.66° 0.18,4.44° 0.17,5.94 ° 0.26,8.45° 0.18,7.28°
c2f-MS-Trans [251] 0.10,4.60° 0.24,9.88° 0.12,12.3° 0.16,5.64° 0.16,4.42° 0.16,6.39° 0.25,7.76° 0.17/7.28°
@@
DSAC [255] 0.02,0.7° 0.03,1.0° 0.02,1.30° 0.03,1.0° 0.05,1.30° 0.05,1.5° 1.90,49.4° 0.30,8.03°
DSAC++[261] 0.02,0.5° 0.02,0.9° 0.01,0.8 0.03,0.7° 0.04,1.1° 0.04,1.1° 0.09,2.6° 0.04,1.10°
$*{}^{*}$ [256] 0.02,1.1° 0.02,1.0° 0.01,1.8° 0.03,1.2° 0.04,1.4° 0.03,1.7° 0.04,1.4° 0.03,1.37°
SANet [43] 0.03/0.9° 0.03/1.1° 0.02/1.5° 0.03/1.0° 0.05/1.3° 0.04/1.4° 0.16/4.6° n.a.
@@
DFNet [277] 0.04,1.48° 0.04,2.16° 0.03,1.82° 0.07,2.01° 0.09,2.26° 0.09,2.42° 0.14,3.31° 0.07,2.21°
Direct-PN+U [276] 0.09,2.77° 0.16,4.87° 0.10,6.64° 0.17,5.04° 0.19,3.59° 0.19,4.79° 0.24,8.52° 0.16,5.17°
LENS [278] 0.03,1.3° 0.10,3.7° 0.07,5.8° 0.07,1.9° 0.08,2.2° 0.09,2.2° 0.14,3.6° 0.08,3.0°
@@
NeRF-Loc [275] 0.02/1.1° 0.02/1.1° 0.01/1.9° 0.02/1.1° 0.03/1.3° 0.03/1.5° 0.03/1.3° n.a.
IX-B How to Select Specific Localization Method?
@@
As shown in the experimental results in Sec. VIII-C, the utilized scene map of MRL usually determines the localization accuracy that a particular kind of MRL can achieve: 1) VPR methods estimate current pose by retrieving historical posed images. Their localization accuracy is limited, but they solely require a 2D image map and can obtain reference images with overlapped view-of-field. As a result, they are often utilized as a coarse step in hierarchical localization framework [23, 24] to acquire co-visible query-reference image pairs. 2) With a slightly better localization accuracy, RPR and APR methods can be employed in some applications that do not necessitate high-precision poses but call for a low-cost and light-weight scene map. 3) VL-MRL and I2P-MRL methods require an offline mapping stage to obtain 3D scene map, i.e., visual landmarks map or point cloud map, so their cost of mapping is significantly higher than previously mentioned methods. With these precise and explicit 3D scene map, these two kind of methods achieve outstanding localization performance under most circumstances. However, when aiming for autonomy in large-scale scenarios, these two kinds of scene map require significant storage space, making them difficult to deploy on resource-limited platform, like intelligent vehicles. 4) HD map is the most light and compact form of scene map, so HD-MRL methods are widely applied in AD task. But HD Map is limited to human-made scenes with semantic map elements, and it is not suitable for natural scenes. 5) Recently developed NN-MRL methods utilize DNN model to represent scenes, which have demonstrated impressive success in small-scale indoor scene and medium-scale urban scenes. However, the efficacy of NN-MRL has yet to be fully evaluated in unbounded scenes.
@@
IX-C What’s Next for Visual Localization?
@@
Developed over several decades, MRL continues to attract growing attention from both academia and industry. We would like to discuss about an interesting and instructive problem: ”What are the future trends of visual localization research?”
@@
End-to-end pipeline: The robustness against environmental changes should be well considered by MRL methods for long-term autonomy. The current query image often has different visual conditions with reference data stored in the scene map. Most of existing solutions rely on image feature algorithms [56, 57] to attain condition-invariant image representation. However, these features often fail to achieve comparable performance to that on feature matching tasks. For instance, the matching accuracy of D2-Net [130] is inferior to SuperPoint [57] on HPatches benchmark [134, 135], but when it comes to localization benchmark [33], D2-Net [130] yields better results than SuperPoint [57]. This discrepancy between feature matching and localization might be attributed to the limitations of human-defined supervision on features. Some recent approaches suggest building the whole localization pipeline into an E2E manner, incorporating feature extraction, matching, and pose estimation, and training such localization models by supervising them on estimated poses [211, 209, 308, 213]. Such pose supervision can facilitate the learning of involved geometry priors in MRL and enhance the global consistency of feature matching. This kind of image feature are specifically learned for MRL, enabling the closure of the gap between feature matching and localization, thus becoming a more reasonable solution for representation in MRL research. It is pertinent to note that HD-MRL is encountered with similar challenges alongside matching-based localization. Since semantic instance-level matching is sparse and difficult, some E2E solutions prefer to make models extract and match high-dimensional features for both images and vectorized map elements, and then estimate poses, leading to improved localization performances [233, 234]. So, we believe E2E localization system will be a promising trend.
@@
Resource-friendly MRL: To achieve real-time ego-pose estimation in large-scale scenarios during long-term exploration, we must take into account the storage demands of scene map and the running efficiency of localization algorithms. For example, high-level autonomous vehicles in unbounded driving areas generally depend on vectorized HD map, rather than visual landmarks map and point cloud map that are commonly used in robotic applications within confined scenes [8, 9]. In the field of robotic localization, several recommendations are available to lighten scene map, e.g., SceneSqueezer [185], NeuMap [28], and HyperMap [201], which significantly decrease the map size while ensuring localization accuracy. There is no doubt that MRL will prefer a lightweight and compact scene map in the future. Aiming at high efficiency, MRL methods are commonly built in coarse-to-fine framework where VPR is applied to coarsely estimate poses [23, 24]. Recent advances enable us to employ NN-MRL solutions as a coarse step, such as APR in LATITUDE [6]. Additionally, improving the efficiency of key components in pipelines is also essential. ALIKE [95] efficiently extracts local features while LightGlue [100] achieves high-speed matching. Efforts to reduce computational cost of MRL methods should be encouraged, as this will stimulate the proposal of more practical MRL solutions for resource-limited platforms.
@@
MRL with new kind of map: In recent years, numerous MRL methods that use new types of scene map have been created. For example, MeshLoc [184] employs a dense mesh-based map rather than visual landmarks map and achieves superior performance. Similarly, NN-MRL methods utilize implicit neural network-based scene map as a substitute for traditional explicit 3D structure-based scene map. Proposing new representation formats of scene map is still an area of research that requires further exploration in MRL. The ideal scene map should be light-weight, compact, and easy to build and deploy. Moreover, it ought to provide comprehensive information for autonomy. The reference data for localization in the scene map should be sufficiently robust against potential changes in real-world scene. Current scene map is challenging to accomplish all of these criteria, so we believe it is necessary to develop new kind of scene map for MRL, which will bring about impressive improvement to MRL researches.
@@
X Conclusions
In this survey, we formulate MRL method as a interaction procedure between query image and scene map where poses are estimated, and then we systemically review MRL methods based on the representation format of utilized scene map, that is, MRL using geo-tagged frame map, visual landmark map, point cloud map, vectorized HD map, and learnt implicit map. Each kind of MRL methods and the related component of them are fully reviewed. Also, we provide a review about evaluation of MRL methods, and draw some conclusions based on the evaluation results of typical algorithms. We provide some opening problem in this field and give our personal opinions. Finally, as a continuous contribution to the community, we list the reviewed papers on the website so that researches can readily find the best-matched MRL method based on their interests.